ii) Calculate the concentration of methanol produced by the reaction in (b), with the


following reactant concentrations:


CO: 2.13 x 10-3 mol L-1


H2: 1.4 x 10-3 mol L-1



1
Expert's answer
2022-02-14T17:06:05-0500

Solution:

CM(CO) = 2.13×10-3 M

CM(H2) = 1.40×10-3 M


Balanced chemical equation:

CO(g) + 2H2(g) ⇌ CH3OH(g)

ICE Table:



Therefore,

[CH3OH] = 0.70×10-3 M = 0.70×10-3 mol L-1


Answer: The concentration of methanol (CH3OH) is 0.70×10-3 mol L-1